OpenStudy (anonymous):

I have another question just like my other one but i dont know how to do this one, Find the solution to each system by the addition (elimination) method....4s-5t=18;6s+10t=-8

OpenStudy (anonymous):

s = 2, t = -2

OpenStudy (anonymous):

ok so how did you do that please explain?

OpenStudy (anonymous):

multiply ur first eqn by 2 and then add it to the second one..ull eliminate the s variable

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so like this 2(4s)-2(5t)=2(18), 8s-10t=36, 6s+10t=-8 8s-10t=36 14s=28 s=14/28=.5

OpenStudy (anonymous):

s=28/14 =2

OpenStudy (anonymous):

right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thats the approach...make the coefficient of one variable the same by multiplying it with a constant..and then eliminate it by adding or subtracting

OpenStudy (anonymous):

ok